2x+3y=8
-x+y=-4
Please help me with linear system

Answer by checkley71(8403)

-X+Y=-4 OR Y=X-4 NOW SUBSTITUTE (X-4) FOR Y IN THE OTHER EQUATION & SOLVE FOR X
2X+3(X-4)=8
2X+3X-12=8
5X=8+12
5X=20
X=20/5
X=4 ANSWER.
-4+Y=-4
Y=-4+4
Y=0 ANSWER.
PROOF
2*4+3*0=8
8=8